Activity 2.1: Review on Finding for the Mean and Variance of the

Population.

Instruction: Do what is being asked.

1. Complete the table by supplying the population mean and population

variance. Round off your answers to hundredths place.


2. What formulas do you use to find the mean and variance of the above given

population


Population Population Mean Population

Variance


1) 1, 2, 3, 4, 5

2) 6, 10, 9, 12, 4, 5

3) 12, 2, 5, 7, 11, 8, 4

4) 14, 10, 15, 14, 17, 19, 20

5) 20, 22, 24, 18, 14, 25, 26, 28


1
Expert's answer
2021-08-18T16:21:20-0400

Solution:

2: Formulae used are: Mean "=\dfrac{\Sigma x_i}{n}" , Variance "=\dfrac{\Sigma (x_i-\bar x)^2}{n}"

1:

1) 1, 2, 3, 4, 5

n = 5

Mean "=\dfrac{1+2+3+4+5}{5}=3"

Variance "=\dfrac{10}{5}=2"

2) 6, 10, 9, 12, 4, 5

n = 6

Mean "=7.6666667"

Variance "=\dfrac{49.333333}{6}=8.22"

3) 12, 2, 5, 7, 11, 8, 4

n = 7

Mean "=7"

Variance "=\dfrac{80}{7}=11.43"

4) 14, 10, 15, 14, 17, 19, 20

n = 7

Mean "=15.571429"

Variance "=\dfrac{69.714286}{7}=9.96"

5) 20, 22, 24, 18, 14, 25, 26, 28

n = 8

Mean "=22.125"

Variance "=\dfrac{148.875}{8}=18.61"